# How does Labeling Error Impact Contrastive Learning? A Perspective from Data Dimensionality Reduction

### v1

@@ Appendix B Lemmas
Assume the set of augmented data $\mathcal{D}_{aug}$ is finite. Let $f^{*}\in\arg\min_{f:\mathcal{D}\rightarrow\mathbb{R}^{k}}$ be a minimizer of the population spectral contrastive loss $\mathcal{L}_{spe}(f)$. Then, there exists a linear head $W^{*}$ with norm $\|W^{*}\|_{F}\leq 1/(1-\lambda_{k})$ such that
$\underset{\tilde{x}\in\tilde{\mathcal{D}},x\in p(\cdot|\tilde{x})}{\mathrm{Pr}}\left[g_{f^{*},W^{*}}(x)\neq y_{\tilde{x}}\right]\leq\frac{4\alpha}{\lambda_{k+1}}+8\alpha.$
@@ Appendix C Proofs of main results
where $V\left(f(x)\right)=\mathbb{E}_{(x,x^{+})\in X^{+}}\left[\left\|f(x)-\mu_{y_{x}}\right\|^{2}\right]$, $V^{-}\left(f(x)\right)=\mathbb{E}_{(x,x^{+})\in X^{-}}\left[\left\|f(x^{+})-\mu_{y_{x}}\right\|^{2}\right]$, $V(f(x^{-}))=V(f|f\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})=\mathbb{E}_{\{f|f\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x}\}}\left[\left\|f-\mu_{y_{f}}\right\|^{2}\right]$ are the intra-class variance of the representations of true positive augmented samples, the variance of false positive augmented samples, and the intra-class variance of negative and true positive augmented samples, respectively.
Proof of Theorem 4.2: Assume positive augmented samples are $x,x^{+}$ and $M$ negative samples are $x_{i}^{-},i=1,...,M$ ($x_{i}^{-}$ belongs to any class in the $K$ classes). We let $\mu_{y_{x}}=\mathbb{E}_{\{x|y=y_{x}\}}\left[f(x)\right]$. Under our assumptions, the labels of $x$ and $x^{+}$ may be different from each other. In other words, data augmentation may change the semantics of the original samples due to the intrinsic randomness. Thus, we denote the sets of true positive sample pair and false positive sample pair as $X^{+}=\left\{(x,x^{+})|y_{x}=y_{x^{+}}\right\}$ and $X^{-}=\left\{(x,x^{+})|y_{x}\neq y_{x^{+}}\right\}$, respectively, and let $X=X^{+}+X^{-}$ be the set of all positive sample pairs $(x,x^{+})$. Then, we have the following lower bounds of the InfoNCE loss
@@
$-\sqrt{V(f(x))}-\sqrt{V^{-}(f(x))}-\frac{1}{2}V(f|f\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})-\mathcal{O}\left(M^{-\frac{1}{2}}\right)-\log\left(\frac{M+1}{K}\right)$
$=$ $\mathcal{L}_{InfoNCE}(f)-\sqrt{V(f(x))}-\sqrt{V^{-}(f(x))}-\frac{1}{2}V(f|f\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})-\mathcal{O}\left(M^{-\frac{1}{2}}\right)-\log\left(\frac{M+1}{K}\right),$
@@
$\geq$ $-\sqrt{V\left(f(x)\right)}-\sqrt{V^{-}\left(f(x)\right)}-\frac{1}{2}V(f|f\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})-\mathcal{O}\left(M^{-\frac{1}{2}}\right).$
@@
$-\frac{1}{2}V(f|f\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})-\mathcal{O}\left(M^{-\frac{1}{2}}\right)-\log\left(\frac{M+1}{K}\right)$
$\geq$ $\mathcal{L}_{InfoNCE}(f)-\epsilon_{q^{*}}^{2}-\epsilon_{q}^{2}-\frac{1}{2}V(f|f\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})-\mathcal{O}\left(M^{-\frac{1}{2}}\right)-\log\left(\frac{M+1}{K}\right).$
@@
Let Assumption 3.1 hold. For the empirical optimal encoder $f^{*}$ taking the truncated SVD with hyper-parameter $q\in[m]$ on the original dataset $\bar{\mathcal{D}}$, there exists a linear head $W$ with norm $\|W^{*}\|_{F}\leq 1/(1-\lambda_{k,q})$ such that $\mathcal{E}(f^{*},W^{*})\leq\frac{4\alpha_{q}}{\lambda_{k+1,q}}+8\alpha_{q}$, where $k$ denotes the dimension of embedding space.
@@
$\underset{\tilde{x}\in\tilde{\mathcal{D}},x\in p(\cdot|\tilde{x})}{\mathrm{Pr}}\left[g_{f^{*},W^{*}}(x)\neq y_{\tilde{x}}\right]\leq\frac{4\alpha}{\lambda_{k+1}}+8\alpha.$
We rewrite the definition $\mathcal{E}(f,W)=\underset{\tilde{x}\in\tilde{\mathcal{D}}}{\mathrm{Pr}}\left[g_{f,W}(\tilde{x})\neq y_{\tilde{x}}\right]$ as $\underset{\tilde{x}\in\tilde{\mathcal{D}},x\in p(\tilde{x}|\tilde{x})}{\mathrm{Pr}}\left[g_{f,W}(x)\neq y_{\tilde{x}}\right]$. Thus, after taking the truncated SVD with hyper-parameter $q\in[m]$ on the labeled dataset $\tilde{\mathcal{D}}$,
$\mathcal{E}(f^{*},W^{*})=\underset{\tilde{x}\in\tilde{\mathcal{D}},x\in p(\tilde{x}|\tilde{x})}{\mathrm{Pr}}\left[g_{f^{*},W^{*}}(x)\neq y_{\tilde{x}}\right]\leq\underset{\tilde{x}\in\tilde{\mathcal{D}},x\in p(\cdot|\tilde{x})}{\mathrm{Pr}}\left[g_{f^{*},W^{*}}(x)\neq y_{\tilde{x}}\right]\leq\frac{4\alpha_{q}}{\lambda_{k+1,q}}+8\alpha_{q}.$

### v2

@@ Appendix B Lemmas
Assume the set of augmented data $\mathcal{D}_{aug}$ is finite. Let $f^{*}\in\arg\min_{f:\mathcal{D}_{aug}\rightarrow\mathbb{R}^{k}}$ be a minimizer of the population spectral contrastive loss $\mathcal{L}_{spe}(f)$ with $k\in\mathcal{Z}^{+}$. Then, there exists a linear head $W^{*}\in\mathbb{R}^{k\times K}$ with norm $\|W^{*}\|_{F}\leq 1/(1-\lambda_{k})$ such that
$\underset{\bar{x}\sim\mathcal{P},x\sim p(\cdot|\bar{x})}{\mathrm{Pr}}\left[g_{f^{*},W^{*}}(x)\neq y_{\bar{x}}\right]\leq\frac{4\alpha}{\lambda_{k+1}}+8\alpha.$
@@ Appendix C Proofs of main results
where $V\left(f(x)\right)=\mathbb{E}_{(x,x^{+})\in X^{+}}\left[\left\|f(x)-\mu_{y_{x}}\right\|^{2}\right]$, $V^{-}\left(f(x)\right)=\mathbb{E}_{(x,x^{+})\in X^{-}}\left[\left\|f(x^{+})-\mu_{y_{x}}\right\|^{2}\right]$, $V(f(x^{-}))=V(z|z\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})=\mathbb{E}_{\{z|z\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x}\}}\left[\left\|z-\mu_{y_{z}}\right\|^{2}\right]$ are the intra-class variance of the representations for true positive augmented samples, the variance for false positive augmented samples, and the intra-class variance for negative and true positive augmented samples, respectively.
Proof of Theorem 4.2: Assume positive augmented samples are $x,x^{+}$ and $M$ negative samples are $x_{i}^{-},i=1,...,M$ ($x_{i}^{-}$ belongs to any class in the $K$ classes). We let $\mu_{y_{x}}=\mathbb{E}_{\{x|y=y_{x}\}}\left[f(x)\right]$. Under our assumptions, the labels of $x$ and $x^{+}$ may be different from each other. In other words, data augmentation may change the semantics of the original samples due to the intrinsic randomness. Thus, we denote the sets of true positive sample pair and false positive sample pair as $X^{+}=\left\{(x,x^{+})|y_{x}=y_{x^{+}}\right\}$ and $X^{-}=\left\{(x,x^{+})|y_{x}\neq y_{x^{+}}\right\}$, respectively. Then, we have the following lower bounds of the InfoNCE loss
@@
$-\sqrt{V(f(x))}-\sqrt{V^{-}(f(x))}-\frac{1}{2}V(z|z\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})-\mathcal{O}\left(M^{-\frac{1}{2}}\right)-\log\left(\frac{M+1}{K}\right)$
$=$ $\mathcal{L}_{InfoNCE}(f)-\sqrt{V(f(x))}-\sqrt{V^{-}(f(x))}-\frac{1}{2}V(z|z\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})-\mathcal{O}\left(M^{-\frac{1}{2}}\right)-\log\left(\frac{M+1}{K}\right),$
@@
$\geq$ $-\sqrt{V\left(f(x)\right)}-\sqrt{V^{-}\left(f(x)\right)}-\frac{1}{2}V(z|z\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})-\mathcal{O}\left(M^{-\frac{1}{2}}\right).$
@@
$-\frac{1}{2}V(z|z\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})-\mathcal{O}\left(M^{-\frac{1}{2}}\right)-\log\left(\frac{M+1}{K}\right)$
$\geq$ $\mathcal{L}_{InfoNCE}(f)-\epsilon_{q^{*}}^{2}-\epsilon_{q}^{2}-\frac{1}{2}V(z|z\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})-\mathcal{O}\left(M^{-\frac{1}{2}}\right)-\log\left(\frac{M+1}{K}\right).$
@@
Let Assumption 3.1 hold. For the empirical optimal encoder $f^{*}$ taking the truncated SVD with hyper-parameter $q$ on unlabeled original dataset, there exists a linear head $W^{*}\in\mathbb{R}^{k\times K}$ with norm $\|W^{*}\|_{F}\leq 1/(1-\lambda_{k,q})$ such that
@@
$\underset{\bar{x}\sim\mathcal{P},x\sim p(\cdot|\bar{x})}{\mathrm{Pr}}\left[g_{f^{*},W^{*}}(x)\neq y_{\bar{x}}\right]\leq\frac{4\alpha}{\lambda_{k+1}}+8\alpha.$
We rewrite the definition $\mathcal{E}(f,W)=\underset{\bar{x}\sim\mathcal{P}}{\mathrm{Pr}}\left[g_{f,W}(\bar{x})\neq y_{\bar{x}}\right]$ as $\underset{\bar{x}\sim\mathcal{P},x\sim p(\bar{x}|\bar{x})}{\mathrm{Pr}}\left[g_{f,W}(x)\neq y_{\bar{x}}\right]$. Thus, after taking the truncated SVD with hyper-parameter $q\in[m]$ on unlabeled sample, the following bound holds
$\mathcal{E}(f^{*},W^{*})=\underset{\bar{x}\sim\mathcal{P},x\sim p(\bar{x}|\bar{x})}{\mathrm{Pr}}\left[g_{f^{*},W^{*}}(x)\neq y_{\bar{x}}\right]\leq\underset{\bar{x}\sim\mathcal{P},x\sim p(\cdot|\bar{x})}{\mathrm{Pr}}\left[g_{f^{*},W^{*}}(x)\neq y_{\bar{x}}\right]\leq\frac{4\alpha_{q}}{\lambda_{k+1,q}}+8\alpha_{q}.$